# Sparse Feature Circuits: Discovering and Editing Interpretable Causal Graphs in Language Models

### v2

@@ A.3 Practical considerations
where ${\mathbf{x}}^{d}$ the copy of ${\mathbf{x}}$ which is downstream of $f_{i}$ in the computation graph, and ${\mathbf{x}}^{u}$ is the copy which is upstream of $f_{i}$. To fix this, we apply a stop gradient to ${\bm{\epsilon}}({\mathbf{x}})$ so that ${\mathbf{x}}^{d}=\hat{{\mathbf{x}}}+\text{stopgrad}({\mathbf{x}}^{u}-\hat{{\mathbf{x}}})$.
@@
Done naively, computing the quantities in (5) and (6) when ${\mathbf{d}}$ or ${\mathbf{m}}\in{\mathbb{R}}^{d_{\text{model}}}$ are SAE errors would take $O(d_{\text{model}})$ backwards passes. Fortunately, one can use the following trick: when $A$ is a constant $1\times n$ matrix, ${\mathbf{x}}\in{\mathbb{R}}^{m}$, and ${\mathbf{y}}={\mathbf{y}}({\mathbf{x}})\in{\mathbb{R}}^{n}$ is a function of ${\mathbf{x}}$, we have
@@
where the right-hand side is a $1\times m$ Jacobian which can be computed with a single backward pass. Thus we can compute (5) with only two backwards passes by first computing $\left.\nabla_{\mathbf{d}}m\right|_{{\mathbf{d}}_{\text{clean}}}$ and then computing $\nabla_{\mathbf{u}}\left(\left.\nabla_{\mathbf{d}}m\right|_{{\mathbf{d}}_{\text{clean}}}\right)$ with another backwards pass, where the second $\left.\nabla_{\mathbf{d}}m\right|_{{\mathbf{d}}_{\text{clean}}}$ is treated as a constant (e.g., by detaching it in Pytorch). A similar trick can be applied for (6).
@@ Appendix B Details on Sparse Autoencoders
B.1 Architecture
Following Bricken et al. (2023), our SAEs are one-layer MLPs with a tied pre-encoder bias. In more detail, our SAEs have parameters
@@ B.1 Architecture
${\mathbf{f}}=\begin{bmatrix}f_{1}({\mathbf{x}})&\dots&f_{d_{\text{SAE}}}({\mathbf{x}})\end{bmatrix}=W_{E}({\mathbf{x}}-{\mathbf{b}}_{D})+{\mathbf{b}}_{E}$
and reconstructions via
@@ B.2 Training
Following Nanda (2023), we cache activations from $10000$ tokens in a buffer and randomly sample batches of size $42^{1}$ 4 for training our SAE. When the buffer is half-depleted, we replenish it with fresh tokens from The Pile. We train for $120000$ steps, resulting in a total of about 2 billion training tokens.
@@
B.3 Evaluation
Here we report on various easy-to-quantify metrics of SAE quality. Note that these metrics leave out important qualitative properties of these SAEs, such as the interpretability of their features (App. D). Our metrics are:
@@ B.3 Evaluation
Appendix C Quality of Linear Approximations of Indirect Effects
Appendix D Human Interpretability Ratings for Sparse Features
Activation type Interpretability
Dense (random) 32.6
Dense (agreement) 30.2
Dense (BiB) 36.0
Sparse (random) 52.8
Sparse (agreement) 62.3
Sparse (BiB) 81.5
Table 7: Human interpretability ratings for dense (neuron) vs. sparse (autoencoder) features. We present mean interpretability scores across features on a 0–100 scale. We show scores for features that were either uniformly sampled (random), the top 30 by $\hat{\text{IE}}$ from the subject–verb agreement across RC task (agreement; §3.3), or the top 30 by $\hat{\text{IE}}$ for the Bias in Bios task (BiB; §4).
We asked human crowdworkers to rate the interpretability of random features, random neurons, features from our feature circuits, and neurons from our neuron circuits on a 0–100 scale (Table 7). Crowdworkers rate sparse features as significantly more interpretable than neurons, with features that participate in our circuits also being more interpretable than randomly sampled features.
See Figures 9 and 10 for examples of the human annotator interface. Humans were presented with the tokens on which the feature activated most strongly, followed by the tokens whose probabilities were most affected in Pythia-70M when the feature was ablated. This is followed by a series of example contexts in which the feature activated on some subset of tokens, where feature activations are shown in varying shades of blue (darker shades indicate higher activations). On the same page below the contexts, we ask annotators to write a textual description of the feature, and rate both its interpretability and its semantic complexity on 0–100 scales.
@@ E.1 Classifier training
Here we describe how we train a classifier on Pythia-70M for the Bias in Bios (BiB) task of §4. To mimic a realistic application setting, we search over hyperparameters to train high-performing baseline and oracle classifiers (using the ambiguous and balanced datasets, respectively). Hyperparameters were not selected for strong Shift performance.
The inputs to our classifier are residual stream activations from the end of the penultimate layer of Pythia-70M. [^5] The implication of this is that we freeze the LM and only update the parameters of the classifier. We mean-pool over (non-padding) tokens from the context In preliminary experiments, we obtained slightly worse baseline and oracle performance when instead extracting representations over only the final token. We also obtained slightly worse performance when training the classifier on activations from the final layer of Pythia-70M.
We then fit a linear probe to these representations with logistic regression using the AdamW optimizer (Loshchilov & Hutter, 2017) with a learning rate of $0.01$ for a single epoch. When retraining after performing Shift, we tune only this linear probe—not the full model.
For unclear reasons, we were unable to fit a probe with greater-than-chance accuracy when applying logistic regression to representations extracted from the final layer; this is why we used penultimate layer representations above.
@@ E.2 Implementation for Concept Bottleneck Probing
Appendix F Feature Circuits
F.1 Subject–Verb Agreement
Here, we present the full agreement circuits for various syntactic agreement structures, where we annotate all features and do not collapse similar features into the same nodes. All circuits here are discovered with $T_{N}=0.1$ and $T_{E}=0.01$ unless otherwise noted. In each circuit, sparse features are shown in rectangles, whereas causally relevant error terms not yet captured by our SAEs are shown in triangles. Nodes shaded in darker colors have stronger effects on the target metric $m$. Blue nodes and edges are those which have positive indirect effects (i.e., are useful for performing the task correctly), whereas red nodes and edges are those which have counterproductive effects on $m$ (i.e., cause the model to consistently predict incorrect answers).
First, we present agreement across a relative clause (Figure 11). The model appears to detect the subject’s grammatical number at the subject position. One position later, features detect the presence of relative pronouns (the start of the distractor clause). Finally, at the last token of the relative clause, the attention moves the subject information to the last position, where it assists in predicting the correct verb inflection.
The circuit for agreement across a prepositional phrase (Figure 12) looks remarkably similar to agreement across a relative clause; these two circuits share over 85% of their features, and many of the same features are used for detecting both prepositions and relative clauses.
For simple agreement (Figure 13, discovered with $T_{N}=0.2$ and $T_{E}=0.02$), many of the same features that were implicated in noun number detection and verb number prediction in the previous circuits also appear here. The model detects the subject’s number at the subject position in early layers. In later layers, these noun number detectors become inputs to verb number promoters, which activate on anything predictive of particular verb inflections.
Finally, the circuit for agreement within a relative clause (Figure 13, discovered with $T_{N}=0.2$ and $T_{E}=0.02$) appears to have the same structure as that for simple agreement: subject number detectors in early layers, followed by verb number promoters in later layers.
Figure 11: The full annotated feature circuit for agreement across a relative clause. The model detects the subject’s number at the subject position. Other features detect relative pronouns (the start of the distractor clause). Finally, at the last token of the RC, the attention moves the subject information to the last position, where it assists in predicting the correct verb inflection.
Figure 12: The full annotated feature circuit for agreement across a prepositional phrase. The model detects the subject’s number at the subject position. Other features detect prepositional phrases (the start of the distractor clause). Finally, at the last token of the RC, the attention moves the subject information to the last position, where it assists in predicting the correct verb inflection.
Figure 13: The full annotated feature circuit for simple agreement. The model detects the subject’s number at the subject position in early layers. In later layers, these are inputs to features which activate on anything predictive of particular verb inflections.
Figure 14: The full annotated feature circuit for agreement within a relative clause. The model detects the subject’s number at the subject (within the RC)’s position in early layers. In later layers, these features are inputs to features which activate on anything predictive of particular verb inflections.
F.2 Bias in Bios Circuit
Here, we present the full annotated circuit discovered for the Bias in Bios classifier (described in §4 and App. E). The circuit was discovered using $T_{N}=0.1$ and $T_{E}=0.01$. We observe that the circuit (Figure 15) contains many nodes which simply detect the presence of gendered pronouns or gendered names. A few features attend to profession information, including one which activates on words related to nursing, and another which activates on passages relating to science and academia.
F.3 Cluster Circuits
Here, we present full annotated circuits discovered for automatically discovered behaviors (described in App. H). First, we present the circuit for incrementing number sequences (Figure 16), discovered with $T_{N}=0.4$ and $T_{E}=0.04$. We note that this circuit includes many features which perform either succession (Gould et al., 2023) or induction (Olsson et al., 2022). The succession features in the layer 3 attention seem to be general; they increment many different numbers and letters (as in Figure 5). The induction features are sensitive only to specific tokens: for example, contexts of the form “$x$3…$x$3”, where “3” is a literal. These compose to form specific successor features in layer 5: the most strongly-activating layer 5 residual feature specifically increments “3” to “4” given induction-like lists, where each list item is preceded by the same string (e.g., “Chapter 1…Chapter 2…Chapter 3…Chapter”).
Appendix G Sample Features
G.1 Sparse Features
G.2 Neurons
Appendix H Discovering LM Behaviors with Clustering
@@ H.1 Filtering Tokens
H.2 Caching Model-internal Information
@@
  • Gradient w.r.t. model parameters: As in Michaud et al. (2023), we also experiment with using gradients of the loss w.r.t. model parameters, but with some modifications. We describe this method in more detail in §H.3 below.
H.3 Hyperparameters and Implementation Details

### v3

@@ A.3 Practical considerations
where ${\mathbf{x}}^{d}$ is the copy of ${\mathbf{x}}$ downstream of $f_{i}$ in the computation graph, and ${\mathbf{x}}^{u}$ is the copy upstream of $f_{i}$. To fix this, we apply a stop gradient to ${\bm{\epsilon}}({\mathbf{x}})$ so that ${\mathbf{x}}^{d}=\hat{{\mathbf{x}}}+\text{stopgrad}({\mathbf{x}}^{u}-\hat{{\mathbf{x}}})$.
@@
Done naively, computing the quantity in (5) when ${\mathbf{d}}$ is an SAE errors would take $O(d_{\text{model}})$ backwards passes. Fortunately, one can use the following trick: when $A$ is a constant $1\times n$ matrix, ${\mathbf{x}}\in{\mathbb{R}}^{m}$, and ${\mathbf{y}}={\mathbf{y}}({\mathbf{x}})\in{\mathbb{R}}^{n}$ is a function of ${\mathbf{x}}$, we have
@@
where the right-hand side is a $1\times m$ Jacobian which can be computed with a single backward pass. Thus we can compute (5) with only two backwards passes by first computing $\left.\nabla_{\mathbf{d}}m\right|_{{\mathbf{d}}_{\text{clean}}}$ and then computing $\nabla_{\mathbf{u}}\left(\left.\nabla_{\mathbf{d}}m\right|_{{\mathbf{d}}_{\text{clean}}}\right)$ with another backwards pass, where the second $\left.\nabla_{\mathbf{d}}m\right|_{{\mathbf{d}}_{\text{clean}}}$ is treated as a constant (e.g., by detaching it in Pytorch).
@@ Appendix B Details on Sparse Autoencoders
B.1 Pythia-70M Sparse Autoencoders
B.1.1 Architecture
Following Bricken et al. (2023), our SAEs for Pythia-70M are one-layer MLPs with a tied pre-encoder bias. In more detail, our SAEs have parameters
@@ B.1.1 Architecture
${\mathbf{f}}=\begin{bmatrix}f_{1}({\mathbf{x}})&\dots&f_{d_{\text{SAE}}}({\mathbf{x}})\end{bmatrix}=\mathrm{ReLU}(W_{E}({\mathbf{x}}-{\mathbf{b}}_{D})+{\mathbf{b}}_{E})$
with the ReLU nonlinearity applied coordinatewise and reconstructions via
@@ B.1.2 Training
Following Nanda (2023), we cache activations from $10000$ contexts in a buffer and randomly sample batches of size $2^{14}$ for training our SAE. When the buffer is half-depleted, we replenish it with fresh tokens from The Pile. We train for $120000$ steps, resulting in a total of about 2 billion training tokens.
@@
B.1.3 Evaluation
Here we report on various easy-to-quantify metrics of SAE quality. Note that these metrics leave out important qualitative properties of these SAEs, such as the interpretability of their features (App. F). Our metrics are:
@@ B.1.3 Evaluation
B.2 Gemma-2-2B Sparse Autoencoders
For Gemma-2-2B, we use the Gemma Scope SAEs released by Lieberum et al. (2024), which are based on the Jump-ReLU architecture proposed by Rajamanoharan et al. (2024b). We use the SAEs of width 16384. There exist SAEs for the attention, MLP, and residual vectors for each of the 26 layers in the model. However, the attention and MLP SAEs are trained at different positions than in Pythia: attention SAEs are trained on the input to the out projection, and MLP SAEs are trained on the output of the LayerNorm following the MLP. The embedding SAEs are experimental and have a dictionary size of only 4000, so we do not use them in our experiments.
There exist multiple SAEs for every submodule. The primary difference between them is their average L0 norm. [^3] In other words, the average number of features active for a given token. Neuronpedia (Lin & Bloom, 2023) uses the SAEs with the L0 norm closest to 100; we do the same.
One of the primary technical challenges in using the Gemma Scope SAEs is the existence of BOS features. These are features that are active primarily or only on BOS tokens, and whose top logits are generally not informative. These features are difficult to interpret, but can have high indirect effects on the model’s logits. As we cannot interpret them, we exclude them from annotation and from the SHIFT analysis (i.e., we do not ablate them). We also exclude them when running the feature skyline in §4.
Appendix C Feature Circuits
C.1 Subject-Verb Agreement
Here, we present the full agreement circuits for various syntactic agreement structures, with researcher-provided annotations for features. We chose thresholds manually in order to keep the number of nodes to annotate manageable while still displaying the full range of feature types for a given task.
In each circuit, sparse features are shown in rectangles, whereas causally relevant error terms not yet captured by our SAEs are shown in triangles. Nodes shaded in darker colors have stronger effects on the target metric $m$. Blue nodes and edges are those which have positive indirect effects (i.e., are useful for performing the task correctly), whereas red nodes and edges are those which have counterproductive effects on $m$ (i.e., cause the model to consistently predict incorrect answers).
First, we present agreement across a relative clause. Pythia (Figure 7) and Gemma (Figure 8) both appear to detect the subject’s grammatical number at the subject position. One position later, features detect the presence of relative pronouns (the start of the distractor clause). Finally, at the last token of the relative clause, the attention moves the subject information to the last position, where it assists in predicting the correct verb inflection. Gemma 2 additionally leverages noun phrase (NP) number tracking features, which are active at all positions for NPs of a given number (except on distractor phrases of opposite number). We present an example of an NP number tracker feature in Figure 15.
The circuits for agreement across a prepositional phrase (Figures 9 and 10) look remarkably similar to agreement across a relative clause; for both Pythia and Gemma, these two circuits share over 85% of their features, and many of the same features are used for detecting both prepositions and relative clauses.
For simple agreement (Figures 11 and 12), many of the same features that were implicated in noun number detection and verb number prediction in the previous circuits also appear here. The models detect the subject’s number at the subject position in early layers. In later layers, these noun number detectors become inputs to verb number promoters, which activate on anything predictive of particular verb inflections.
The circuits for agreement within a relative clause (Figures 13 and 14) appear to have the same structure as that for simple agreement: subject number detectors in early layers, followed by verb number promoters in later layers.
Figure 7: The feature circuit for agreement across a relative clause in Pythia-70M, computed using $T_{N}=0.1$ and $T_{E}=0.01$. The model detects the subject’s number at the subject position. Other features detect relative pronouns (the start of the distractor clause). Finally, at the last token of the RC, the attention moves the subject information to the last position, where it assists in predicting the correct verb inflection.
Figure 8: The feature circuit for agreement across a relative clause in Gemma-2-2B, computed using $T_{N}=0.073$ and $T_{E}=0.007$.
Figure 9: The feature circuit for agreement across a prepositional phrase in Pythia-70M, computed using $T_{N}=0.1$ and $T_{E}=0.01$. The model detects the subject’s number at the subject position. Other features detect prepositional phrases (the start of the distractor clause). Finally, at the last token of the RC, the attention moves the subject information to the last position, where it assists in predicting the correct verb inflection.
Figure 10: The feature circuit for agreement across a prepositional phrase in Gemma-2-2B, computed using $T_{N}=0.5$ and $T_{E}=0.05$. Note that we show the circuit beginning in layer 13, as our circuit discovery implicated only one node in earlier layers.
Figure 11: The feature circuit for simple agreement in Pythia-70M, computed using $T_{N}=0.2$ and $T_{E}=0.02$. The model detects the subject’s number at the subject position in early layers. In later layers, these are inputs to features which activate on anything predictive of particular verb inflections.
Figure 12: The feature circuit for simple agreement in Gemma-2-2B, computed using $T_{N}=0.5$ and $T_{E}=0.05$.
Figure 13: The feature circuit for agreement within a relative clause in Pythia-70M, computed with $T_{N}=0.2$ and $T_{E}=0.02$. The model detects the subject’s number at the subject (within the RC)’s position in early layers. In later layers, these features are inputs to features which activate on anything predictive of particular verb inflections.
Figure 14: The feature circuit for agreement within a relative clause for Gemma-2-2B, computed with $T_{N}=0.5$ and $T_{E}=0.05$.
Figure 15: An example sparse feature for agreement across a relative clause in Gemma 2 (resid_12/13561). This feature activates on tokens in noun phrases where the noun head is plural, but not on singular distractor phrases within the plural NP. This feature carries the number of the subject across positions, so we term it an “NP number tracker”.
C.2 Bias in Bios Circuit
Here, we present the full annotated circuit discovered for the Bias in Bios classifier trained on Pythia-70M (described in §4 and App. E). The circuit was discovered using $T_{N}=0.1$ and $T_{E}=0.01$. We observe that the circuit (Figure 16) contains many nodes which simply detect the presence of gendered pronouns or gendered names. A few features attend to profession information, including one which activates on words related to nursing, and another which activates on passages relating to science and academia.
C.3 Cluster Circuits
Here, we present full annotated circuits discovered for automatically discovered behaviors (described in App. G). First, we present the circuit for incrementing number sequences (Figure 17), discovered with $T_{N}=0.4$ and $T_{E}=0.04$. We note that this circuit includes many features which perform either succession (Gould et al., 2023) or induction (Olsson et al., 2022). The succession features in the layer 3 attention seem to be general; they increment many different numbers and letters (as in Figure 5). The induction features are sensitive only to specific tokens: for example, contexts of the form “$x$3…$x$3”, where “3” is a literal. These compose to form specific successor features in layer 5: the most strongly-activating layer 5 residual feature specifically increments “3” to “4” given induction-like lists, where each list item is preceded by the same string (e.g., “Chapter 1…Chapter 2…Chapter 3…Chapter”).
Appendix D Sample Features
D.1 Sparse Features
D.2 Neurons
@@ E.1 Classifier training
Here we describe how we train linear classification heads on Pythia-70M and Gemma-2-2B the Bias in Bios (BiB) task of §4.
Given a model $M$ and choice $\ell$ of layer, we mean-pool over (non-padding) tokens all layer $\ell$ residual stream activations from $M$; we then train a linear classification head via logistic regression, using the AdamW optimizer (Loshchilov & Hutter, 2017) and learning rate $0.01$ for one epoch on this dataset of activations. The activations and labels for this logistic regression are collected from the ambiguous set for the baseline classifier and from the balanced set for the oracle classifier.
To mimick a realistic application setting, we tune the choice $\ell$ of layer for the baseline probe’s accuracy on (a test split of) the ambiguous set. For Pythia, this recommends using the penultimate layer $\ell=4$. For Gemma, there is a wide range of equally performant layers. Thus—for this one choice only—we make use of the balanced set to compute how the baseline probe generalizes; we select the layer $\ell=22$ for which the baseline probe generalizes worst. We make this choice to set up a testbed where there is the most space for improvement. We emphasize that we never tune hyperparameters for the performance balanced set of Shift, as using the balanced set is forbidden by the problem statement.
When retraining after performing Shift, we retrain only the linear classification head, not the full model.
@@ E.2 Implementation for Concept Bottleneck Probing
Appendix F Human Interpretability Ratings for Sparse Features
Activation type Interpretability
Dense (random) 32.6
Dense (agreement) 30.2
Dense (BiB) 36.0
Sparse (random) 52.8
Sparse (agreement) 62.3
Sparse (BiB) 81.5
Table 7: Human interpretability ratings for dense (neuron) vs. sparse (autoencoder) features. We present mean interpretability scores across features on a 0–100 scale. We show scores for features that were either uniformly sampled (random), the top 30 by $\hat{\text{IE}}$ from the subject-verb agreement across RC task (agreement; §3.3), or the top 30 by $\hat{\text{IE}}$ for the Bias in Bios task (BiB; §4).
Given our trained Pythia-70M sparse autoencoders, we asked human crowdworkers to rate the interpretability of random features, random neurons, features from our feature circuits, and neurons from our neuron circuits on a 0–100 scale (Table 7). Crowdworkers rate sparse features as significantly more interpretable than neurons, with features that participate in our circuits also being more interpretable than randomly sampled features.
See Figures 23 and 24 for examples of the human annotator interface. Humans were presented with the tokens on which the feature activated most strongly, followed by the tokens whose probabilities were most affected in Pythia-70M when the feature was ablated. This is followed by a series of example contexts in which the feature activated on some subset of tokens, where feature activations are shown in varying shades of blue (darker shades indicate higher activations). On the same page below the contexts, we ask annotators to write a textual description of the feature, and rate both its interpretability and its semantic complexity on 0–100 scales.
Results of human evaluations for the Gemma Scope SAEs are described in Lieberum et al. (2024).
Appendix G Discovering LM Behaviors with Clustering
@@ G.1 Filtering Tokens
G.2 Caching Model-internal Information
@@
  • Gradient w.r.t. model parameters: As in Michaud et al. (2023), we also experiment with using gradients of the loss w.r.t. model parameters, but with some modifications. We describe this method in more detail in §G.3 below.
G.3 Hyperparameters and Implementation Details
@@
Appendix H Quality of Linear Approximations of Indirect Effects